OpenStudy (anonymous):

@whpalmer4 solve 3^(2x)=108

OpenStudy (anonymous):

The left side can be simplified to 3^2 * 3^x, which yields 9*3^x. We divide both sides by nine to get: 3^x = 9 Now we have a simple equation we can solve using logarithms. Do you know how to solve this?

OpenStudy (anonymous):

\[\log_{3} 108=2x\]

OpenStudy (whpalmer4):

\[3^{2x} = 108\]\[\ln 3^{2x} = \ln 108\]\[2x\ln 3 = \ln 108\]\[x = \frac{\ln 108}{2\ln 3}\approx 2.13\]

OpenStudy (anonymous):

Okay, or do it the inefficient way. I am sad now.

OpenStudy (whpalmer4):

I'm sad that you think 108/9 = 9!

OpenStudy (anonymous):

wait is in ln or log?

OpenStudy (anonymous):

I'm really good at math, I promise. But 3^x = 12 is still a much easier equation to solve.

OpenStudy (anonymous):

Arithmetic gets to me sometimes. :C Anyway, ln is the natural logarithm. Log is used typically in reference to the common logarithm unless otherwise specified.

OpenStudy (anonymous):

i came up with 2.5719 ... not 2.13

OpenStudy (anonymous):

@murdocjax check your answer.

OpenStudy (whpalmer4):

Well, I wrote out steps for the benefit of the chap who asked me for help reviewing for his test. Normally, I would just write down \[x = \frac{\ln 108}{2 \ln 3}\]without any need to factor out 3^2 and divide both sides. Being able to look at the problem and write the answer immediately makes me think my method is satisfactory, especially when it doesn't cause me to do things that sometimes trip me up, but there are certainly many ways to solve such problems, and everyone should find the way that works best for their needs.

OpenStudy (whpalmer4):

\[3^{2*2.5719} \approx 284.588\]

OpenStudy (anonymous):

yes the long way helps to understand the problem

OpenStudy (anonymous):

wait now i'm confused. what's the final answer?

OpenStudy (whpalmer4):

Also, \[3^{2x} \ne 9*3^x\] \[3^{2x} = 3^x*3^x\]

OpenStudy (whpalmer4):

(well, okay, \(3^{2x} = 9*3^x\) if \(x=2\)) @murdocjax my original answer is correct.

OpenStudy (anonymous):

I don't know, man. I have a really weird intuition of exponents, so I break things apart a lot.

OpenStudy (anonymous):

the 2.13

OpenStudy (whpalmer4):

key word might be "break" :-) Here's a plot of \(3^{2x}\) and \(9*3^x\) which I think conclusively demonstrates that they are not equal for just about every value of interest here.

OpenStudy (whpalmer4):

@murdocjax you can do the problem with logs of any base, and still get the same answer. For example, if we do it with \(\log_{10}\): \[x = \frac{\log_{10} 108}{2\log_{10}3} = \frac{2.03342}{2*0.477121} = 2.13093\] Do it with \(\log_3\): \[x = \frac{\log_3 108}{2\log_3 3} = \frac{4.26186}{2*1} = 2.13093\] Do it with \(\ln\): \[x = \frac{\ln 108}{2 \ln 3} = \frac{ 4.68213}{2*1.09861} = 2.13093\]

OpenStudy (anonymous):

@whpalmer, I think you may want to check your order of operations for those graphs.

OpenStudy (anonymous):

I think I meant 9^x. Yeah, I meant that. I'm messing up a lot. Oh well.

OpenStudy (whpalmer4):

uh huh. http://www.wolframalpha.com/input/?i=3%5E%282x%29%2C+9*3%5Ex+from+x%3D1+to+5 Look, you are asserting that \[3^2*3^x = 3^{2x}\] Let's try out some values. \[x=0\]\[3^2*3^0 = 3^{2*0}\]\[9*1 = 1\]Uh, no. \[x=1\] \[3^2*3^1 = 3^{2*1}\]\[9*3 = 3^2\]\[27=9\]Uh, no.

OpenStudy (anonymous):

"I think I meant 9^x. Yeah, I meant that. I'm messing up a lot. Oh well."

OpenStudy (anonymous):

And Wolfram confirms me on that one. http://www.wolframalpha.com/input/?i=3%5E%282*x%29

OpenStudy (whpalmer4):

Yes, \[((x^m)^n) = x^{mn}\] By now the chap has had time to memorize log 2 and log 3 in a number of bases, which is all he would need to evaluate that result as a decimal. \[108 = 2^2*3^3\]\[\log 108 = 2\log 2 + 3\log 3\]